2x(y+z^2)p+y(2y+z^2)q=z^2
And prove thatyz(z^2+2z-2y)=x^2
1
Expert's answer
2019-04-29T10:22:54-0400

1)

"2x(y+z^2)\partial z/\partial x+y(2y+z^2)\partial z/\partial y=z^2"

"dx\/(2x(y+z^2))=dy\/(y(2y+z^2))=dz\/z^2"

"ydx\/(2xy^2+2xyz^2)=xdy\/(2xy^2+xyz^2))=xydz\/(xyz^2)"

"(ydx-xdy-xydz)\/(2xy^2+2xyz^2-2xy^2-xyz^2-xyz^2)=(ydx-xdy-xydz)\/0"

"ydx-xdy-xydz=0"

"dx\/x-dy\/y-dz=0"

"lnx-lny-z=c"

"z=ln(x\/y)+C"

2)

"yz(z^2+2z-2y)=x^2"

"yz^3+2yz^2-2y^2z=x^2"

Differentiate respect to x:


"3yz^2p+4yzp-2y^2p=2x"

"py(3z^2+4z-2y)=2x"

Differentiate respect to y:


"z^3+3z^2yq+2z^2+4yzq-4yz-2y^2q=0"

"qy(3z^2+4z-2y)+z^3+2z^2-4yz=0"

So we have:


"3z^2+4z-2y=2x\/(py)"

"2xq\/p+z^3+2z^2-4yz=0"

Since


"p=1\/x"

"q=-1\/y"

then:


"-2x^2\/y+z^3+2z^2-4yz=0"

"-2x^2+y(z^3+2z^2-4yz)=0"

"yz(z^2+2z-4y)=2x^2"

So:


"(z^2+2z-4y)\/(z^2+2z-2y)=2"

"z^2+2z-4y=2z^2+4z-4y"

"z^2+2z=0"

So we get that the statement can be proved if


"z=0"

or


"z=-2"